### v1

# An alternative GPU acceleration for a pseudopotential plane-waves density functional theory code with applications to metallic systems

Abstract
We present an alternative GPU acceleration for plane waves pseudopotentials electronic structure codes designed for systems that have small unit cells but require a large number of k points to sample the Brillouin zone as happens, for instance, in metals. We discuss the diagonalization of the Kohn and Sham equations and the solution of the linear system derived in density functional perturbation theory. Both problems take advantage from a rewriting of the routine that applies the Hamiltonian to the Bloch wave-functions to work simultaneously (in parallel on the GPU threads) on the wave-functions with different wave-vectors k, as many as allowed by the GPU memory. Our implementation is written in CUDA Fortran and makes extensive use of kernel routines that run on the GPU (GLOBAL routines) or can be called from inside the GPU threads (DEVICE routines). We compare our method with the CPUs only calculation and with the approach currently implemented in Quantum ESPRESSO that uses GPU accelerated libraries for the FFT and for the linear algebra tasks such as the matrix-matrix multiplications as well as OpenACC directives for loop parallelization. We show in a realistic example that our method can give a significant improvement in the cases for which it has been designed.
keywords:
Electronic structure; Lattice dynamics; Metals; GPU; Density functional perturbation theory.
journal: Computer Physics Communications
1 Introduction
Density functional theory [1] (DFT) and the availability of more and more powerful computers has made the study of material properties from first-principles a well established reality. Several tools have been refined over the years to solve the one electron Kohn and Sham equations that derive from DFT, [2] the most widespread being based on a plane waves basis and pseudopotentials. Well tested, freely available [3, 4, 5] or commercial [6, 7] packages implement the theory and allow the calculation of material properties.
In the last ten years, high performance computers aiming to reach the exaflops ($10^{18}$ floating point operations per second) switched to a hybrid technology in which the graphic processing units (GPUs) support the central processing units (CPUs) in the floating point operations. In theory, the GPUs can deliver one or two orders of magnitude more flops than the CPUs themselves, and to harness this power many electronic structure groups are modifying their codes to run on the GPUs[8, 9, 10, 11, 12, 13]
Extensions of common programming languages such as C/C++ or Fortran have provided commands to allocate data on the GPU, to move data from the CPU to the GPU and vice versa, and to perform calculations on these data with the GPU. CUDA Fortran commands, [14] declarations, and compiler directives and OpenACC compiler directives [15] are two of the most commonly used Fortran extensions. Recently also applications based on openMP started to appear in the literature. [16, 17] Actually OpenACC and OpenMP have the additional benefit of being transferable to GPUs architectures of different vendors such as NVIDIA, AMD, or Intel and sometimes are preferred to CUDA Fortran that is limited to NVIDIA GPUs.
So far, in several plane-waves pseudopotentials and quantum chemistry codes, the GPUs have been exploited by allocating the variables on the GPU and by substituting the calls to linear algebra and fast Fourier transform (FFTs) routines with calls to optimized library routines (such as cuBlas[18] cuSolver[19] cuFFT[20] and MAGMA [21]) developed by the GPUs vendors and capable to run on the GPU[22, 9, 16, 23, 24, 25] Sometimes the routines of these libraries have the same names and arguments of the corresponding CPU libraries and it suffices to allocate the variables on the GPU to call the GPU routines with minimal changes to the underlying codes and algorithms. Single loops using variables allocated on the GPUs can also be accelerated by compiler directives.
In Quantum ESPRESSO [3, 4] work on this kind of acceleration started more than ten years ago [26] and has been improved over the years [27] leading to a well tested package. [9] Accelerations of 2X or higher with respect to the CPU are often found in pseudopotential plane waves codes that adopt this approach. However, test systems are usually big supercells with many atoms for which the time spent to make calculations on the GPU is larger than the time needed to transfer data from the CPU to the GPU, while small size systems are left out from these tests. For some applications, metallic systems with small unit cells need tens or hundreds thousands k points to sample the Fermi surface. [28, 29, 30, 31, 32, 33, 34, 35, 36] The calculation of the phonon dispersions of these systems for many geometries as required for thermodynamic calculations is a problem that could take advantage from the new supercomputers, but for small systems we found that the use of the GPUs with the present codes is not always convenient and sometimes it can also slow down the calculation with respect to the CPUs alone.
We have therefore tried to improve the situation and in this paper we present the solution that we have found: an alternative approach to the acceleration of the pseudopotentials plane waves codes that is useful to deal with metallic system when there are many k points. We load on the GPU many wave-functions (i.e. k points), all the available ones if the GPU memory is large enough or as many as possible until there is free GPU memory. Then we make the calculations simultaneously on all these data (application of the Hamiltonian to the wave-functions) with each GPU thread working on a single wave-function or on a part of it. To obtain the precise control of the GPU threads that is needed we wrote a set of kernel functions (called GLOBAL in the CUDA language) that implement the theory and run on the GPU. These kernel functions need to call linear algebra and FFT library functions from inside the GPU threads. Unfortunately, libraries such as cuFFT, cuSolver, or MAGMA, which are called from the CPU and automatically control the number of threads, are not suited for our algorithm. We need functions that can be called from inside the GPU threads (DEVICE functions in the CUDA language). We only find the C++ library cuFFTDx [37] that implements such functionalities, but it does not provide a FORTRAN interface so far. For the moment, we transformed to the DEVICE form the FORTRAN sources of fftpack5.1 and of selected LAPACK routines. Finally, we obtained a code significantly faster than the standard one for small systems with many k points.
We start with a brief introduction of the main equations that are solved in a plane-waves pseudopotential code. We stress in particular the algorithms that are relevant for the following discussion, neglecting the parts that have not changed or that are still calculated on the CPU. We then discuss how, in our method, the different parts of the code have been accelerated on the GPU. Finally, we present a test of our implementation and compare the times required by our approach with those taken by the CPUs only calculations and by the previously available GPU implementation.
2 Theory
The solutions of the Kohn and Sham (KS) equations minimize the DFT total energy. These equations are an eigenvalue problem for norm conserving pseudopotentials, [38] and a generalized eigenvalue problem for ultrasoft [39] or projector-augmented wave (PAW) pseudopotentials. [40, 41] For periodic solids they can be written as:
${H}_{KS}\psi_{{\bf k}\nu}=\varepsilon_{{\bf k}\nu}S\psi_{{\bf k}\nu},$ (1)
where ${\bf k}$ is a wave vector and $\nu$ is a band index. $H_{KS}$ is the Kohn and Sham Hamiltonian and $S$ is the overlap matrix. We are interested in finding the lowest $N_{b}$ (number of bands) eigenvalues and eigenvectors of these equations. The Kohn and Sham Hamiltonian depends itself from a potential that is calculated from the charge density (that also depends on the wavefunctions). It is possible to solve this problem by a self-consistent procedure in which the wavefunctions are first calculated with an approximate potential. Then these wavefunctions are used to recompute the charge density and a new potential. The latter is mixed with the potential of the previous iterations and the procedure is repeated until one reaches self-consistency. At each step of the procedure however one has to diagonalize a fixed Hamiltonian which is progressively improved.
In the standard algorithm the problem is solved sequentially for each ${\bf k}$ vector and the charge density is computed at the end when all wave-functions are available.
2.1 Davidson algorithm
There are several algorithms currently implemented in electronic structure codes to find the eigenvalues and eigenfunctions in Eq. 1, but here we limit the discussion to the Davidson algorithm. [42] In this algorithm an initial set of $N_{b}$ functions $|\phi_{i}^{(n)}\rangle$ are progressively improved by enlarging the set applying $H_{KS}-\varepsilon_{i}S$ and solving the generalized eigenvalue problem $\tilde{H}_{ij}-\varepsilon\tilde{S}_{ij}$ on the basis formed by the original and the newly calculated vectors. A standard software library for numerical linear algebra, such as LAPACK[43] is employed for the diagonalization. The algoritm is the following and has to be repeated for each k point:
  • 1.
    Given $N_{b}$ trial eigenpairs: $\quad\left\{\left|\phi_{i}^{(n)}\right\rangle,\varepsilon_{i}^{(n)}\right\}$ of the reduced Hamiltonian calculate:
    $\tilde{H}_{ij}=\left\langle\phi_{i}^{(n)}\left|H_{KS}\right|\phi_{j}^{(n)}\right\rangle,\quad\tilde{S}_{ij}=\left\langle\phi_{i}^{(n)}|S|\phi_{j}^{(n)}\right\rangle.$ (2)
  • 2.
    Build the correction vectors $\left|\tilde{\phi}_{i}^{(n)}\right\rangle$:
    $\left|\tilde{\phi}_{i}^{(n)}\right\rangle=\left(H_{diag}-\varepsilon_{i}^{(n)}S_{diag}\right)^{-1}\left(H_{KS}-\varepsilon_{i}^{(n)}S\right)\left|\phi_{i}^{(n)}\right\rangle,$ (3)
    where $H_{diag}$ and $S_{diag}$ are the diagonal elements of $H_{KS}$ and $S$ in the plane waves representation.
  • 3.
    Normalize the correction vectors:
    $|\tilde{\phi}_{i}^{(n)}\rangle={|\tilde{\phi}_{i}^{(n)}\rangle\over\sqrt{\langle\tilde{\phi}_{i}^{(n)}|\tilde{\phi}_{i}^{(n)}\rangle}}.$ (4)
  • 4.
    Build an extended reduced Hamiltonian and overlap matrix:
    $\tilde{H}_{ij}=\left\langle\phi_{i}^{(n)}/\tilde{\phi}_{i}^{(n)}\left|H_{KS}\right|\phi_{j}^{(n)}/\tilde{\phi}_{j}^{(n)}\right\rangle,\quad\tilde{S}_{ij}=\left\langle\phi_{i}^{(n)}/\tilde{\phi}_{i}^{(n)}|S|\phi_{j}^{(n)}/\tilde{\phi}_{j}^{(n)}\right\rangle.$ (5)
  • 1.
    Set $N_{base}$ equal to the number of basis vector. Diagonalize the small $N_{base}\times N_{base}$ reduced Hamiltonian to get the new estimate for the lowest $N_{b}$ eigenpairs:
    $(\tilde{H}-\varepsilon\tilde{S})v=0\quad\\ \longrightarrow\quad\left\{\left|\phi_{i}^{(n+1)}\right\rangle,\varepsilon_{i}^{(n+1)}\right\}.$ (6)
  • 2.
    Calculate $|\tilde{\phi_{i}}^{(n+1)}\rangle$ for all $i$ for which $\left|\varepsilon_{i}^{(n+1)}-\varepsilon_{i}^{(n)}\right|>\varepsilon_{th}$ where $\varepsilon_{th}$ is the accuracy required for the eigenvalues and call $N_{nc}$ the number of new vectors.
  • 3.
    If $N_{nc}>0$ repeat with the basis $|\phi_{i}^{(n)}/\tilde{\phi_{i}}^{(n)}/\tilde{\phi_{i}}^{(n+1)}\rangle$ of size $N_{base}+N_{nc}$ and continue with progressively larger basis. When the size of the basis becomes too large for the allocated memory instead of adding $\quad\left\{\left|\tilde{\phi_{i}}^{(n+1)}\right\rangle\right\}$ to the basis, restart with $\quad\left\{\left|\phi_{i}^{(n+1)}\right\rangle,\varepsilon_{i}^{(n+1)}\right\}$. If $N_{nc}=0$ exit with eigenpairs $\quad\left\{\left|\phi_{i}^{(n+1)}\right\rangle,\varepsilon_{i}^{(n+1)}\right\}$
The most time consuming step of this algorithm is the application of the operators $H_{KS}$ and $S$ to the wave-functions as discussed in the next subsection.
2.2 Application of Hamiltonian
The KS Hamiltonian can be written as: [44]
${H_{KS}}\psi_{{\bf k}\nu}=\underbrace{-\frac{1}{2}\nabla^{2}\psi_{{\bf k}\nu}}_{kinetic\ energy}+\underbrace{V_{eff}\psi_{{\bf k}\nu}}_{local\ energy}+\underbrace{V_{NL}\psi_{{\bf k}\nu}}_{non-local\ energy},$ (7)
where the effective potentials is the sum of the local, Hartree, and exchange and correlation potentials:
$V_{eff}=V_{loc}+V_{H}+V_{XC},$ (8)
while the nonlocal pseudopotential is defined in term of the projector functions $\left|\beta^{I}_{m}\right\rangle$ and pseudopotential coefficients $D^{I}_{mn}$[38, 39]
$V_{NL}\left|\psi_{{\bf k}\nu}\right\rangle=\sum_{Imn}{D^{I}_{mn}}{\left|\beta^{I}_{m}\right\rangle}{\left\langle\beta^{I}_{n}\right|\psi_{{\bf k}\nu}\big{\rangle}}.$ (9)
Here $I$ indicates the different atoms in the solid and $m$ and $n$ run on all the $\beta^{I}_{m}$ functions of a given atom.
The overlap matrix can be calculated in a similar way: [39]
$S\left|\psi_{{\bf k}\nu}\right\rangle=\left|\psi_{{\bf k}\nu}\right\rangle+\sum_{mn}q^{I}_{mn}\left|\beta^{I}_{m}\right\rangle\left\langle\beta^{I}_{n}\big{|}\psi_{{\bf k}\nu}\right\rangle,$ (10)
where the coefficients $q^{I}_{mn}$ are defined together with the pseudopotential.
2.2.1 Kinetic energy
The kinetic energy is calculated in reciprocal space. Using the Bloch theorem we write the Bloch wave-functions as:
$\psi_{{\bf k}\nu}({\bf r})=e^{i{\bf k}{\bf r}}u_{{\bf k}\nu}({\bf r})=\frac{1}{\sqrt{V}}\sum_{{\bf G}}C_{{\bf k}+{\bf G}\nu}e^{i({\bf k}+{\bf G}){\bf r}},$ (11)
where $u_{{\bf k}\nu}({\bf r})$ is a lattice periodic function expanded in plane waves (here $V$ is the volume of the solid) and the sum is over the reciprocal lattice vectors contained into a sphere defined by the relationship:
${1\over 2}|{\bf k}+{\bf G}|^{2}<E_{cut},$ (12)
where $E_{cut}$ is the kinetic energy cut-off. Then we have:
$-{1\over 2}\nabla^{2}\psi_{{\bf k}\nu}({\bf r})={1\over\sqrt{V}}\sum_{{\bf G}}C^{\prime}_{{\bf k}+{\bf G}\nu}e^{i({\bf k}+{\bf G}){\bf r}},$ (13)
where:
$C^{\prime}_{{\bf k}+{\bf G}\nu}={1\over 2}|{\bf k}+{\bf G}|^{2}C_{{\bf k}+{\bf G}\nu}.$ (14)
2.2.2 Local potential
The fast Fourier transform (FFT) transforms functions in real space into reciprocal space and the inverse FFT makes the inverse transformation.
From the coefficients $C_{{\bf k}+{\bf G}\nu}$, applying an inverse FFT we obtain the Bloch function in real space (up to a factor $1/\sqrt{V}$):
$C_{{\bf k}+{\bf G}\nu}\xrightarrow{FFT^{-1}}u_{{\bf k}\nu}({\bf r})=\sum_{{\bf G}}C_{{\bf k}+{\bf G}\nu}e^{i{\bf G}{\bf r}}.$ (15)
The effective potential is applied in real space as:
$u^{\prime}_{{\bf k}\nu}({\bf r})=V_{eff}({\bf r})u_{{\bf k}\nu}({\bf r}),$ (16)
and a final FFT computes the plane wave expansion of $u^{\prime}_{{\bf k}\nu}({\bf r})$:
$u^{\prime}_{{\bf k}\nu}({\bf r})\xrightarrow{{FFT}}C^{\prime}_{{\bf k}+{\bf G}\nu}={1\over N_{\bf r}}\sum_{{\bf r}}u^{\prime}_{{\bf k}\nu}({\bf r})e^{-i{\bf G}{\bf r}},$ (17)
where $N_{\bf r}$ is the number of points of the FFT grid (see below).
The actual calculation of $V_{eff}$ requires the calculation of the charge density in terms of the wave-functions $\psi_{{\bf k}\nu}$. However since we have not modified this part of the calculation we do not discuss it in detail. We assume only to have a function $V_{eff}$ defined in the points of the FFT grid ${\bf r}$.
2.2.3 Non local pseudopotential and overlap matrix
The application of the non local potential needs three matrix-matrix multiplications:
$\lambda^{I}_{n{\bf k}\nu}=\left\langle\beta^{I}_{n}\big{|}\psi_{{\bf k}\nu}\right\rangle=\sum_{\bf G}\beta^{I}_{n}({\bf k}+{\bf G})^{*}C_{{\bf k}+{\bf G}\nu},$ (18)
$\gamma^{I}_{m{\bf k}\nu}=\sum_{n}{D^{I}_{mn}}\lambda^{I}_{n{\bf k}\nu},$ (19)
$V_{NL}\left|\psi_{{\bf k}\nu}\right\rangle=\sum_{Im}\gamma^{I}_{m{\bf k}\nu}\left|\beta^{I}_{m}\right\rangle.$ (20)
Similarly, the application of the $S$ matrix is:
$\delta^{I}_{m{\bf k}\nu}=\sum_{n}{q^{I}_{mn}}\lambda^{I}_{n{\bf k}\nu},$ (21)
$S\left|\psi_{{\bf k}\nu}\right\rangle=\left|\psi_{{\bf k}\nu}\right\rangle+\sum_{Im}\delta^{I}_{m{\bf k}\nu}\left|\beta^{I}_{m}\right\rangle.$ (22)
where $\lambda^{I}_{n{\bf k}\nu}$ are those calculated in Eq. 18.
2.3 Density functional perturbation theory
The phonon frequencies and displacement modes are obtained by diagonalization of the dynamical matrix:
$\omega_{\mathbf{q}}^{2}\mathbf{u}_{s\alpha}(\mathbf{q})=\sum_{s^{\prime}\beta}D_{s\alpha s^{\prime}\beta}(\mathbf{q})\mathbf{u}_{s^{\prime}\beta}(\mathbf{q}),$ (23)
where $D_{s\alpha s^{\prime}\beta}(\mathbf{q})$ is the dynamical matrix:
$D_{s\alpha s^{\prime}\beta}(\mathbf{q})=\frac{1}{\sqrt{M_{s}M_{s^{\prime}}}}\sum_{\nu}\frac{\partial^{2}E_{tot}}{\partial\mathbf{u}_{\mu s\alpha}\partial\mathbf{u}_{\nu s^{\prime}\beta}}e^{i\mathbf{q}\left(\mathbf{R}_{\nu}-\mathbf{R}_{\mu}\right)},$ (24)
where $E_{tot}$ is the DFT total energy, ${\bf q}$ is a wave vector in the Brillouin zone (BZ), $\mathbf{R}_{\mu}$ are the Bravais lattice vectors, $M_{s}$ are the atomic masses, and $\mathbf{u}_{\mu s\alpha}$ are the atomic displacements.
The second derivative of the DFT total energy can be written in terms of the change of the wave-functions due to a phonon perturbation projected on the conduction band. These functions are the solutions of a linear system: [45, 46]
$\left[H_{KS}^{{\bf k}+{\bf q}}+\alpha Q^{{\bf k}+{\bf q}}-\varepsilon_{{\bf k}\nu}S\right]P_{c}^{{\bf k}+{\bf q}}\frac{\partial u_{{\bf k}\nu}(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}=-P_{c}^{{\bf k}+{\bf q}}\left[\frac{\partial V_{KS}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}-\varepsilon_{{\bf k},\nu}\frac{\partial S}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}\right]u_{{\bf k}\nu}(\mathbf{r}),$ (25)
where $P_{c}^{{\bf k}+{\bf q}}$ is the projector in the conduction band and $\frac{\partial V_{KS}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}=\frac{\partial V_{\text{loc }}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}+\frac{\partial V_{H}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}+\frac{\partial V_{\text{xc }}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}+\frac{\partial V_{\text{NL }}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}$. The change of the Hartree and exchange and correlation potential are:
$\frac{\partial V_{H}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}$ $=\int\frac{e^{i{\bf q}({\bf r}^{\prime}-{\bf r})}}{\left|\mathbf{r}-\mathbf{r}^{\prime}\right|}\frac{\partial\rho\left(\mathbf{r}^{\prime}\right)}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}d^{3}r^{\prime},$ (26)
$\frac{\partial V_{xc}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}$ $=\frac{dV_{xc}}{d\rho}\frac{\partial\rho(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})},$
and depend self-consistently on the charge density induced by the perturbation:
$\frac{\partial\rho(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}=4\sum_{{\bf k}\nu}\left[u_{{\bf k}\nu}^{*}(\mathbf{r})P_{c}^{{\bf k}+{\bf q}}\frac{\partial u_{{\bf k}\nu}(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}\right]+\Delta_{{\bf u}_{s^{\prime}\beta}({\bf q})}({\bf r}),$ (27)
where the last term represents the change of the augmentation charge calculated in the ultrasoft and PAW case but not accelerated in the present work. [46] $Q^{{\bf k}+{\bf q}}$ is a projection in the valence manifold, [45] while the change of the nonlocal pseudopotential is described in more detail in the given references (see for instance Ref. [46]).
$\alpha Q^{{\bf k}+{\bf q}}$ can be written in the form:
$\alpha Q^{{\bf k}+{\bf q}}=\alpha\sum_{\mu}S|u_{{\bf k}+{\bf q}\mu}\rangle\langle u_{{\bf k}+{\bf q}\mu}|S,$ (28)
and its application to a set of wave functions $|x_{{\bf k}+{\bf q}\nu j}\rangle$ (here $j$ indicates the different perturbations ${s^{\prime}\beta}$) can be calculated easily using the fact that $S|x_{{\bf k}+{\bf q}\nu j}\rangle$ is already known from the routine that applies $H^{{\bf k}+{\bf q}}_{KS}$ and $S$. We have a first matrix-matrix multiplication:
$\mu_{{\bf k}+{\bf q}j\mu\nu}=\langle u_{{\bf k}+{\bf q}\mu}|S|x_{{\bf k}+{\bf q}\nu j}\rangle\alpha,$ (29)
then a second one:
$|y_{{\bf k}+{\bf q}\nu j}\rangle=\sum_{\mu}|u_{{\bf k}+{\bf q}\mu}\rangle\mu_{{\bf k}+{\bf q}j\mu\nu},$ (30)
and finally we must apply $S$ to the vectors $|y_{{\bf k}+{\bf q}\nu j}\rangle$ and we have:
$\alpha Q^{{\bf k}+{\bf q}}|x_{{\bf k}+{\bf q}\nu j}\rangle=S|y_{{\bf k}+{\bf q}\nu j}\rangle=|y_{{\bf k}+{\bf q}\nu j}\rangle+\sum_{Imn}q^{I}_{mn}|\beta^{I}_{m}\rangle\langle\beta^{I}_{n}|y_{{\bf k}+{\bf q}\nu j}\rangle,$ (31)
and this requires other three matrix-matrix multiplications as illustrated above.
The self-consistent linear system (Eq. 25) is solved by iterations. From an initial guess of the potentials, of ${\partial V_{H}\over{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}}+{\partial V_{xc}\over{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}}$, the linear system is solved and new induced charge and potentials are obtained. Mixing the latter with the potentials used in the linear system it is possible to reach a self-consistent solution.
The most time consuming step of this process is however the solution of the linear system at fixed ${\partial V_{H}\over{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}}+{\partial V_{xc}\over{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}}$ so we will focus on this step.
2.4 Preconditioned conjugate gradient
The algorithm used for the solution of Eq. 25 with a given right hand side is a preconditioned conjugate-gradient iterative algorithm. [47, 48, 49] Given a starting guess $x$ of the solution of the problem $Ax=b$, we improve it with the following algorithm:
$r$$\leftarrow$ $Ax-b,$ (32)
$d$$\leftarrow$ $M^{-1}r,$ (33)
$\rho$$\leftarrow$ $d^{T}r,$ (34)
$\gamma$$\leftarrow$ ${\rho\over\rho_{old}},$ (35)
$d$$\leftarrow$ $d+\gamma d_{old},$ (36)
$t$$\leftarrow$ $Ad,$ (37)
$\lambda$$\leftarrow$ $-{d^{T}r\over d^{T}t},$ (38)
$x$$\leftarrow$ $x+\lambda d,$ (39)
$r$$\leftarrow$ $r+\lambda t,$ (40)
$d_{old}$$\leftarrow$ $d,$ (41)
$\rho_{old}$$\leftarrow$ $\rho,$ (42)
and iterate from Eq. 33 until the modulus of $\rho$ is smaller than an input threshold. Here the arrows indicate that the variables on the left are substituted with those on the right. $r$ is the negative of the residual vector while $d$ contains minus the preconditioned residual in Eqs. 33 to 36 and minus the search direction from Eq. 36. Eqs. 35 and 36 are executed only from the second iteration onwards. The algorithm requires memory sufficient to save the vectors $r$, $d$, $t$, $d_{old}$ of the same size of the input vector $x$. $\rho$ and $\rho_{old}$, as well as $\gamma$ and $\lambda$, are instead scalars. Moreover, we need two external routines to apply $A$ and $M^{-1}$. The most time consuming step is the application of the matrix $A$ to $d$. In our case $A=H^{{\bf k}+{\bf q}}_{KS}+\alpha Q^{{\bf k}+{\bf q}}-\varepsilon_{{\bf k}\nu}S$ so again the acceleration rests on the routine that applies $H^{{\bf k}+{\bf q}}_{KS}$ and $S$ described above. For the preconditioning the following matrix diagonal in reciprocal space $M_{{\bf G},{\bf G}}=MAX(1.0,{|{\bf k}+{\bf q}+{\bf G}|^{2}\over 2\langle\psi_{{\bf k}+{\bf q}\nu}|-{1\over 2}\nabla^{2}|\psi_{{\bf k}+{\bf q}\nu}\rangle})$ is used and this vector is passed to the routine. The conjugate gradient algorithm is applied to each ${\bf k}$ point and to each $N_{pe}$ perturbations. The $N_{b}$ bands of a given ${\bf k}$ point are optimized together but the different k points and different perturbations are treated in sequence, one after the other. Only one ${\bf q}$ is calculated in each run. This algorithm has been used in the last thirty years in Quantum ESPRESSO to solve the linear system. Similar algorithms, with appropriate modifications, can be used also to minimize the total energy and solve the Kohn and Sham equations. [50]
3 GPU optimization
Figure 1: Algorithms used in the standard approach and in our optimized GPU approach for the diagonalization of the Hamiltonian.
Algorithm 1 CPU and standard GPU diagonalization
for ik = 1,nks do $\triangleright$ nks = #k points per pool
     build $H_{KS}^{\bf k}$ and $S^{\bf k}$
     compute $\varepsilon_{{\bf k}\nu}$ and $\psi_{{\bf k}\nu}$ by Davidson ($N_{b}$)
end for
Algorithm 2 Optimized GPU diagonalization
for ikb = 1,nkblock do $\triangleright$ nkblock = #k points blocks
     build in parallel $H_{KS}^{\bf k}$ and $S^{\bf k}$ on GPU threads ($N_{k}$)
     compute $\varepsilon_{{\bf k}\nu}$ and $\psi_{{\bf k}\nu}$ by Davidson on GPU threads ($N_{b}\times N_{k}$)
end for
Figure 2: Flowchart of the routine that applies $H^{\bf k}_{KS}$ and $S^{\bf k}$ to the wavefunctions. Close to each routine we write the number of threads that are used to run it on the GPU.
Quantum ESPRESSO has several levels of parallelization on the CPU. It is possible to divide the k-points in groups (called pools) and assign each group to a set of cores. These cores may be further divided in groups, with each group dealing with a set of bands (bands parallelization) and finally each group of cores dealing with a set of bands can further divide the reciprocal lattice vectors (G) and work only on a subset of these (G vectors parallelization). It is at this point that one can introduce the GPU acceleration. The standard method to use the GPU consists into allocating variables on the device memory (the GPU) and to call from HOST routines developed by NVidia that perform linear algebra (cuBLAS) operations or FFTs (cuFFT) on the data allocated on the device. It is also possible to add compiler directives to run loops in parallel on the GPU without changing the code.
The standard Davidson algorithm works sequentially on each ${\bf k}$ point of a pool and solves for all the bands (calling the routines that apply $H_{KS}$ and $S$ for a subset of bands, if bands parallelization is used). Therefore the number of times in which the GPU memory is loaded increases linearly with the number of ${\bf k}$ points. When the size of the problem is small, it can happen that the library matrix-matrix multiplications and FFTs routines cannot exploit all the capacity of the GPU because they have too few data to work on. As a result a GPU calculation might become even slower than a CPUs only calculation. Parallelizing on the ${\bf G}$ vectors just reduces further the size of the data allocated on the GPU for each k point and does not help in this case. Moreover, presently the GPU acceleration does not work well if many CPUs use the same GPU, so we use as many CPUs as GPUs.
Our strategy for accelerating the code on the GPU is illustrated schematically in Fig. 1. We put on the GPU memory as many wave-functions (i.e. k points) as possible in a block of $N_{k}$ k points and run simultaneously on all these k points the operations of the Davidson algorithm needed to diagonalize the Hamiltonian. Each pool of CPUs cores works on its set of k points as assigned by the pool parallelization of Quantum ESPRESSO and only these k points are divided in blocks for the GPU acceleration. Bands paralellization and G vector parallelization presently are not supported by our approach. The main GPU optimization has been performed on the routine that applies $H_{KS}$ and $S$ to the wave-functions $\psi_{{\bf k}\nu}$, but some acceleration has been obtained also carrying out the operations of the Davidson algorithm in parallel on many k points. In our approach, the routine that applies $H_{KS}$ and $S$ is a HOST routine (i.e. a routine running on the CPU) that receives as input $C_{{\bf k}+{\bf G}\nu}$ for $N_{k}$ k points, and gives as output the coefficients $C^{\prime}_{{\bf k}+{\bf G}\nu}$ and $C^{\prime\prime}_{{\bf k}+{\bf G}\nu}$ of the plane waves expansion of $H_{KS}\psi_{{\bf k}\nu}$ and of $S\psi_{{\bf k}\nu}$. This routine calls in sequence several GLOBAL routines (that is routines that run on the GPU and for which we can specify how many GPU threads run in parallel). The sequence of routines and the formula that they implement is illustrated in Fig. 2. The first computes the kinetic energy and runs $N_{pw}\times N_{k}\times N_{b}$ threads each one dealing with a ${\bf G}$ vector of one k point and of one band (here $N_{pw}$ is the number of ${\bf G}$ vectors used to expand the wave-functions). A second routine computes the scalar product in Eq. 18 and runs $N_{kb}\times N_{k}\times N_{b}$ threads, where $N_{kb}$ is the total number of projectors $|\beta^{I}_{m}\rangle$. The latter are loaded on the GPU for all the $N_{k}$ points before calling the Davidson algorithm. Another GLOBAL routine computes Eqs. 19 and 21 and runs $N_{k}\times N_{b}$ threads, while the sum over $n$ is made inside the routine. A routine copies $C_{{\bf k}+{\bf G}\nu}$ in $C^{\prime\prime}_{{\bf k}+{\bf G}\nu}$ and this is made in parallel running $N_{pw}\times N_{k}\times N_{b}$ threads. This is the first term of the application of $S$ to the wave functions. A routine sets to zero the FFT grid running $N_{\bf r}\times N_{k}\times N_{b}$ threads and another one sets the non zero elements of this grid running $N_{k}\times N_{b}$ threads, each one dealing with all the $N_{\bf r}$ grid points for one ${\bf k}$ point and one band. Then a set of three routines applies an inverse FFT to the wave-functions as detailed below, and a routine applies $V_{eff}$ running $N_{\bf r}\times N_{k}\times N_{b}$ threads. Another set of three routines applies the FFT to return to reciprocal space and a routine collects the results from the grid and adds them to $C^{\prime}_{{\bf k}+{\bf G}\nu}$. This is made in parallel on the GPU running $N_{k}\times N_{b}$ threads. Finally Eq. 20 is calculated by a routine that runs $N_{pw}\times N_{k}\times N_{b}$ threads and adds the result to $C^{\prime}_{{\bf k}+{\bf G}\nu}$. In the ultrasoft or PAW PPs case, the same routine calculates also the second term in the right hand side of Eq. 22 and adds it to $C^{\prime\prime}_{{\bf k}+{\bf G}\nu}$.
In this algorithm, $N_{k}$ must be carefully chosen and depends on the amount of GPU memory and on the size of the FFT grid. $N_{k}$ is mainly limited by the necessity to allocate on the GPU $N_{k}\times N_{b}$ FFT grids to apply, in parallel, the local potential to the Bloch functions. The allocation of this memory is done by the HOST routine that implements the Davidson algorithm.
We have also optimized some parts of the Davidson algorithm. The standard routine has been generalized introducing several loops on the $N_{k}$ k points and part of these loops have been transformed into kernel routines (GLOBAL routines) that perform the calculation in parallel using $N_{k}\times N_{b}$ threads. We have accelerated only the loops that took a significant amount of time. The other loops call the linear algebra cuBlas routines as in the standard approach.
Figure 3: Algorithms used in the standard phonon code and in our optimized GPU approach for solving the linear system that gives the perturbed wavefunctions.
Algorithm 3 CPU and standard GPU phonon algorithm
for ik = 1,nks do
     build $H_{KS}^{{\bf k}+{\bf q}}$ and $S^{{\bf k}+{\bf q}}$
     for ipert = 1,npe do $\triangleright$ npe = #perturbations
         build $P_{c}^{{\bf k}+{\bf q}}\frac{\partial V_{KS}}{\partial u_{s\alpha}({\bf q})}u_{{\bf k}v}$ ($N_{b}$)
         compute $\frac{\partial u_{{\bf k}v}}{\partial u_{s\alpha}({\bf q})}$ by conjugate gradient (CG) ($N_{b}$)
     end for
end for
Algorithm 4 Optimized GPU phonon algorithm
for ikb = 1,nkblock do $\triangleright$ $N_{k}$ = #k points per block
     build in parallel $H^{{\bf k}+{\bf q}}_{KS}$ and $S^{{\bf k}+{\bf q}}$ on GPU threads ($N_{k}$)
     build in parallel $P_{c}^{{\bf k}+{\bf q}}\frac{\partial V_{KS}}{\partial u_{s\alpha}({\bf q})}u_{{\bf k}v}$ on GPU threads ($N_{b}\times N_{k}\times N_{pe}$)
     compute $\frac{\partial u_{{\bf k}v}}{\partial u_{s\alpha}({\bf q})}$ by CG on GPU threads ($N_{b}\times N_{k}\times N_{pe}$)
end for
The acceleration of the phonon code instead has been carried out essentially on the algorithm that solves the linear system in Eq. 25. We proceed as in the Davidson algorithm (see the scheme in Fig. 3). However, in the phonon case the calculation of the induced charge density (Eq. 27) requires two FFT grids per band, one to contain $u_{{\bf k}\nu}^{*}(\mathbf{r})$ and one to contain $P_{c}^{{\bf k}+{\bf q}}\frac{\partial u_{{\bf k}\nu}(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}$ so usually we use $N_{k}$ smaller than in the Davidson algorithm. The GPU optimization of the preconditioned conjugate gradient algorithm starts by allocating the COMPLEX vectors $g$, $d$, $d_{old}$, and $t$ on the GPU. For each variable $N_{k}\times N_{b}\times N_{pe}$ arrays are allocated. This memory is much larger than the one of the standard algorithm that requires only $N_{b}$ copies of each variable, but this space is allocated only on the GPU. The algorithm is then divided in loops over the $N_{k}$ k points, the $N_{pe}$ perturbations, and the $N_{b}$ bands. Loop one executes Eqs. 33 and Eqs. 34, loop two executes Eq. 36, and loop three computes $d^{T}r$ and $d^{T}t$ that appear in the numerator and denominator of Eq. 38. Finally loop four computes Eqs. 39, 40, and 41. Each loop is transformed into routine with the GLOBAL attribute that runs $N_{k}\times N_{b}\times N_{pe}$ threads, each one computing one perturbation to one band of one k point. Since each thread executes only a scalar product or an operation of the type $x\leftarrow x+\lambda d$ we have programmed these routines in CUDA Fortran without calling any other library routine. The array themselves instead are not split and each thread works on all the G vectors of each wave-function. All the other steps of the algorithm involve only scalar operations that are performed by the CPU.
Eq. 32 and Eq. 37 require an external routine to apply $A$ to the vectors $x$ (or $d$). For this operator we use the same routine that applies $H_{KS}$ and $S$ in the Davidson algorithm. The routine works in general for an arbitrary number of wavefunctions so when called from the conjugate gradient algorithm, in parallel on the GPU threads, it deals with the $N_{k}\times N_{pe}$ set of wavefunctions, each one composed by $N_{b}$ bands. We have then written a HOST routine that receives as input the coefficients $C^{\prime}_{{\bf k}+{\bf q}\nu j}$ and $C^{\prime\prime}_{{\bf k}+{\bf q}\nu j}$ of the Fourier transform of $H^{{\bf k}+{\bf q}}_{KS}|x_{{\bf k}+{\bf q}\nu j}\rangle$ and $S|x_{{\bf k}+{\bf q}\nu j}\rangle$ and gives as output the Fourier coefficients of $A|x_{{\bf k}+{\bf q}\nu j}\rangle$. This routine calls a series of GLOBAL routines for which we can control the number GPU threads that run in parallel. The first routine computes Eq. 29 and runs $N_{k}\times N_{pe}\times N_{b}\times N_{b}$ threads. A second routine computes $|a_{{\bf k}+{\bf q}\nu j}\rangle=H_{KS}^{{\bf k}+{\bf q}}|x_{{\bf k}+{\bf q}\nu j}\rangle-\varepsilon_{{\bf k}\nu}S|x_{{\bf k}+{\bf q}\nu j}\rangle$ and runs $N_{k}\times N_{pe}\times N_{b}$ threads. To complete the operator $A$ we have to calculate the operator $\alpha Q^{{\bf k}+{\bf q}}|x_{{\bf k}+{\bf q}\nu j}\rangle$ and we optimized also this part to run in many threads on the GPU in parallel on the ${\bf k}$ vectors, the bands, and the perturbations. This is done by calling another set of GLOBAL routines. The first computes Eq. 30 running on $N_{k}\times N_{pe}\times N_{b}$ threads, another one computes the scalar products $\langle\beta^{I}_{n}|y_{{\bf k}+{\bf q}\nu j}\rangle$ that appear in Eq. 31 and runs $N_{k}\times N_{pe}\times N_{kb}\times N_{b}$ threads, and a third routine calculates Eq. 21 using the scalar products just calculated and runs on $N_{k}\times N_{pe}\times N_{b}$ threads. Finally a GLOBAL routine computes Eq. 31 and adds it to $|a_{{\bf k}+{\bf q}\nu j}\rangle$ running in $N_{k}\times N_{pe}\times N_{b}\times N_{pw}$ threads.
4 Fast Fourier transform
The application of $V_{eff}({\bf r})$ to one Bloch wave-function requires two Fourier transforms. It is convenient to introduce a mesh in reciprocal space:
${\bf G}_{m_{1},m_{2},m_{3}}\equiv m_{1}{\bf b}_{1}+m_{2}{\bf b}_{2}+m_{3}{\bf b}_{3},$ (43)
where ${\bf b}_{1}$, ${\bf b}_{2}$, and ${\bf b}_{3}$ are the principal reciprocal lattice vectors and $m_{1}$, $m_{2}$, and $m_{3}$ are integers, and a mesh in real space:
${\bf r}_{l_{1},l_{2},l_{3}}={l_{1}\over N_{1}}{\bf a}_{1}+{l_{2}\over N_{2}}{\bf a}_{2}+{l_{3}\over N_{3}}{\bf a}_{3},$ (44)
where ${\bf a}_{1}$, ${\bf a}_{2}$, and ${\bf a}_{3}$ are the direct lattice vectors, and $l_{1}$, $l_{2}$, and $l_{3}$ are integers. The integers $N_{1}$, $N_{2}$, and $N_{3}$ define the size of the mesh in real space and, equivalently, the size of the mesh in reciprocal space. They must be sufficiently large so that the vectors ${\bf G}_{m_{1},m_{2},m_{3}}$ contain all the vectors ${\bf G}-{\bf G}^{\prime}$ defined by the basis set.
[…]